OpenStudy (anonymous):

write an equation in standard form of the line that passes through the given point and has the given slope m or that passes through the two given points (answer in form Ax-By=C (3,0),m=-4/7

OpenStudy (gorv):

y=mx+c \[y=\frac{ -4 }{ 7 }*x+c\] put value of x and y to find c \[0=\frac{ -4 }{ 7 } *3+c=\frac{ -12 }{ 7 }+c\]

OpenStudy (gorv):

\[c=\frac{ 12 }{ 7 }\]

OpenStudy (gorv):

\[y=\frac{ -4 }{7}*x+\frac{ 12 }{ 7 }\]

OpenStudy (gorv):

\[7y=-4*x+12\] 4x+7y=12

OpenStudy (gorv):

A=4 B=-7 c=12